OpenStudy (anonymous):

@phi Nathan had an infection, and his doctor wanted him to take penicillin. Because Nathan’s father and paternal grandfather were allergic to penicillin, Nathan has a 75% chance of having the same allergy. The doctor performed a skin test to see whether Nathan would react to it. The test is 98% accurate. What is the probability that Nathan is not allergic to penicillin and the test predicts it? A. 0.005 B. 0.015 C. 0.245 D. 0.735

OpenStudy (phi):

Nathan has 0.75 chance of allergy 0.25 chance of no allergy test is 0.98 correct 0.02 incorrect Nathan is not allergic to penicillin and the test predicts it or in other words Nathan no allergy and test incorrect multiply the probabilities: Pr(no allergy) * Pr(incorrect) = 0.25*0.02 = 0.005

OpenStudy (anonymous):

i get what you did with your method.

OpenStudy (anonymous):

Answer 0.005 was wrong though. Thanks though );

OpenStudy (lyssakat):

Its C!